# MiniDrive: More Efficient Vision-Language Models with Multi-Level 2D Features as Text Tokens for Autonomous Driving

### v1

@@ 1 Introduction
To address these challenges, this paper introduces a novel vision-language model called MiniDrive. Unlike traditional mainstream visual-language models, MiniDrive is not a unified model based on the Transformer architecture. We use the efficient backbone network model UniRepLKNet Ding et al. (2024), which is based on large convolutional kernels, as the vision encoder. We propose the Feature Engineering Mixture of Experts (FE-MoE) and the Dynamic Instruction Adapter (DI-Adapter) to sequentially process visual features and obtain visual tokens before inputting them into the language model. Specifically, UniRepLKNet captures the 2D features of images, and FE-MoE processes multiple 2D features, mapping them into text tokens for input into the language model without requiring stage-wise training for cross-modal fine-grained alignment. Additionally, the DI-Adapter is introduced to enable the mapped visual tokens (i.e., text tokens used as input to the language model) to dynamically adapt to user text instructions, effectively aiding cross-modal understanding between text and images. As shown in Figure 1(a), MiniDrive processes multiple input images along with user instructions to generate natural language responses. It encompasses the most critical capabilities in autonomous driving, including perception, planning, and prediction question-answering abilities. In Figure 1(b), we illustrate that MiniDrive is a lightweight visual-language model with a minimal parameter size, memory footprint, and FLOP count. It can be fully trained with multiple instances on a single RTX 4090 GPU with 24GB of memory. For instance, $\text{MiniDrive}_{224}$ has only 83M parameters and a FLOP count of merely 5.9B, which is significantly lower than current visual-language models used in autonomous driving. In terms of response performance, MiniDrive outperforms a series of previous models in question-answering capabilities. Notably, its response quality exceeds that of models with billions of parameters. Additionally, MiniDrive supports both single and multiple image inputs. In Figure 1(c), MiniDrive outperforms open-source models with 7B parameters and above on the single-image evaluation system CODA-LM Li et al. (2024), approaching the performance of closed-source commercial models.
@@
Here are our main contributions: (1) We develop autonomous driving VLMs—MiniDrive, which address the challenges of efficient deployment and real-time response in VLMs for autonomous driving systems while maintaining excellent performance. The training cost of the model is reduced, and multiple MiniDrive models can be fully trained simultaneously on an RTX 4090 GPU with 24GB of memory. (2) MinDrive is attempting for the first time to utilize a large convolutional kernel architecture as the vision encoder backbone for autonomous driving vision-language models, enabling more efficient and faster extraction of 2D features at different image levels. We propose Feature Engineering Mixture of Experts (FE-MoE), which addresses the challenge of efficiently encoding 2D features from multiple perspectives into text token embeddings, effectively reducing the number of visual feature tokens and minimizing feature redundancy. (3) This paper introduces the Dynamic Instruction Adapter through a residual structure, which addresses the problem of fixed visual tokens for the same image before being input into the language model. The DI-Adapter enables visual features to dynamically adapt to different textual instructions, thereby enhancing cross-modal understanding. (4) We conduct extensive experiments on MiniDrive, achieving state-of-the-art performance compared to autonomous driving VLMs with multi-view image inputs on Drive-LM. Further, we outperform general open-source VLMs (>7B) with single-image inputs on CODA-LM by an average of 13.2 points. We open-source all our resources to foster community development.
@@ 2.2 Autonomous Driving Based on LLMs
While existing work provides significant value and demonstrates strong capabilities for autonomous driving, most models have over a billion parameters. They are largely based on large-scale language models such as GPT-3.5 and LLaMA, and rely on vision encoders built on the ViT architecture, such as CLIP, ViT-g/14, and ViT-B/32. This results in high computational costs, making these models unsuitable for online scenarios. Although there is a trend towards developing lightweight models for autonomous driving, their performance still falls short compared to larger models.

### v3

@@ 1 Introduction
To address these challenges, this paper introduces a novel vision-language model called MiniDrive. Unlike traditional mainstream visual-language models, MiniDrive is not a unified model based on the Transformer architecture. We use the efficient backbone network model UniRepLKNet Ding et al. (2024), which is based on large convolutional kernels, as the vision encoder. We propose the Feature Engineering Mixture of Experts (FE-MoE) and the Dynamic Instruction Adapter (DI-Adapter) to sequentially process visual features and obtain visual tokens before inputting them into the language model. Specifically, UniRepLKNet captures the 2D features of images, and FE-MoE processes multiple 2D features, mapping them into text tokens for input into the language model without requiring stage-wise training for cross-modal fine-grained alignment. Additionally, the DI-Adapter is introduced to enable the mapped visual tokens (i.e., text tokens used as input to the language model) to dynamically adapt to user text instructions, effectively aiding cross-modal understanding between text and images. As shown in Figure 1(a), MiniDrive processes multiple input images along with user instructions to generate natural language responses. It encompasses the most critical capabilities in autonomous driving, including perception, planning, and prediction question-answering abilities. In Figure 1(b), we illustrate that MiniDrive is a lightweight visual-language model with a minimal parameter size, memory footprint, and FLOP count. It can be fully trained with multiple instances on a single RTX 4090 GPU with 24GB of memory. For instance, $\text{MiniDrive}_{224}$ has only 83M parameters and a FLOP count of merely 5.9B, which is significantly lower than current visual-language models used in autonomous driving. In terms of response performance, MiniDrive outperforms a series of previous models in question-answering capabilities. Notably, its response quality exceeds that of models with billions of parameters. Additionally, MiniDrive supports both single and multiple image inputs. In Figure 1(c), MiniDrive outperforms open-source models with 7B parameters and above on the single-image evaluation system CODA-LM Li et al. (2024), approaching the performance of closed-source commercial models. Here are our main contributions:
  • We develop autonomous driving VLMs—MiniDrive, which address the challenges of efficient deployment and real-time response in VLMs for autonomous driving systems while maintaining excellent performance. The training cost of the model is reduced, and multiple MiniDrive models can be fully trained simultaneously on an RTX 4090 GPU with 24GB of memory.
@@
  • MinDrive is attempting for the first time to utilize a large convolutional kernel architecture as the vision encoder backbone for autonomous driving vision-language models, enabling more efficient and faster extraction of 2D features at different image levels. We propose Feature Engineering Mixture of Experts (FE-MoE), which addresses the challenge of efficiently encoding 2D features from multiple perspectives into text token embeddings, effectively reducing the number of visual feature tokens and minimizing feature redundancy.
  • This paper introduces the Dynamic Instruction Adapter through a residual structure, which addresses the problem of fixed visual tokens for the same image before being input into the language model. The DI-Adapter enables visual features to dynamically adapt to different textual instructions, thereby enhancing cross-modal understanding.
  • We conduct extensive experiments on MiniDrive, achieving state-of-the-art performance compared to autonomous driving VLMs with multi-view image inputs on Drive-LM. Further, we outperform general open-source VLMs (7B) with single-image inputs on CODA-LM by an average of 13.2 points. We open-source all our resources to foster community development.
Figure 2: Comparison of the MiniDrive architecture with mainstream architectures. (a) Existing vision-language models primarily use a Transformer-based visual encoder to learn image patches as visual tokens. These visual tokens remain unchanged regardless of the user’s questions. (b) Our architecture employs a more efficient large convolutional kernel as the visual encoder, learning 2D features of the image as visual tokens. These visual tokens change in response to different user questions.
@@ 2.2 Autonomous Driving Based on LLMs
While existing work provides significant value and demonstrates strong capabilities for autonomous driving, most models have over a billion parameters. They are largely based on large-scale language models such as GPT-3.5 and LLaMA, and rely on vision encoders built on the ViT architecture, such as CLIP, ViT-g/14, and ViT-B/32. This results in high computational costs, making these models unsuitable for online scenarios. Although there is a trend towards developing lightweight models for autonomous driving, their performance still falls short compared to larger models. In Figure 2, we summarize the architectures of the current mainstream vision-language models and compare them with the architecture of MiniDrive. Existing vision-language models primarily divide images into several patches using a Transformer-based visual encoder, learning each patch as tokens for the input language model. Additionally, during inference, the visual tokens remain fixed regardless of how the user’s query changes. Our architecture extracts multi-level 2D features from the image using a vision encoder based on large convolutional kernels (LKNet), further extracting and learning these features to map them as tokens for the input language model. CNNs with large convolutional kernels are more efficient and lightweight (Ding et al., 2022, 2024). Meanwhile, changes in the user’s query dynamically alter the visual tokens.